OpenStudy (anonymous):

Having a square bedroom with 5 mtrs as the length of the diagonal, compute for the number of 300mmx 300mm floor tile.. can you answer this question? :) thnx!!~

OpenStudy (anonymous):

If the room is square the diagonal will cut it into 2 right triangles. Each triangle has 2 45 degree angle and a 90 degree angle. Therefore, the diagonal is the hypotenuse of the triangle.

OpenStudy (anonymous):

Do you understand so far?

OpenStudy (anonymous):

With a right triangle you can find the sides of the triangle using trigonometric identities. Let us use sine. \[Sin \theta=\frac{ opposite }{ hypotenuse }\]

OpenStudy (anonymous):

 Sine 45 = x/5 \[x=5\frac{\sqrt{2} }{ 2 }\] Convert the tiles sizes from mm to meter makes it 0.03mmx0.03mm It would take x/0.03 =117.85 300mm x 300mm tiles to cover the width of the room and also the length. round 117.85 to 118 tiles To find the number of tiles 118*118 = 13924 tiles

OpenStudy (anonymous):

Any questions?

OpenStudy (anonymous):

I made a mistake replace 0.03m with 0.3 instead.

OpenStudy (anonymous):

300mm = 0.3m

OpenStudy (anonymous):

you should get 11.79, round up to 12 tiles to cover the width of the room. the length is also 12: therefore it would take 12*12 = 144 tiles

OpenStudy (anonymous):

Sorry about the computation error. Apart from that do you see how i got the answer?

OpenStudy (anonymous):

Yes, It is a reasonable answer...